# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 2.3.3 Appendix: $C^{\infty}$-description of the induced infinitesimal deformation
For any complex submanifold $S^{\prime}$ of $S$, let $(\mathcal{E}_{S^{\prime}},\theta_{\mathcal{E},S^{\prime}})$ denote the restriction of $(\mathcal{E},\theta_{\mathcal{E}})$ to $S^{\prime}\times X$. Let $q_{S^{\prime}}:S^{\prime}\times X\to X$ denote the projection. Let $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S^{\prime}}$ denote the sheaf of $C^{\infty}$-functions $f$ on $S^{\prime}\times X$ such that ¯$\overline{\partial}_{S^{\prime}}f=0$.
@@
Let $\varphi_{U_{0}}$ be the isomorphism $q_{S_{2}}^{\ast}(\mathcal{E})_{|S_{2}\times U_{0}}\simeq\mathcal{E}_{|S_{2}\times U_{0}}$ defined by $\varphi_{U_{0}}(q_{S_{2}}^{\ast}({\boldsymbol{v}}_{0|\{s_{0}\}\times U_{0}}))={\boldsymbol{v}}_{0|SS_{2}\times U_{0}}$. Let $\varphi_{U_{i}}$ be the isomorphism $q_{S_{2}}^{\ast}(\mathcal{E})_{|S_{2}\times U_{i}}\simeq\mathcal{E}_{|S_{2}\times U_{i}}$ defined by
@@
Proof   Because ¯$\overline{\partial}_{q_{S_{0}}^{\ast}(E)}(B^{1,0})+\bigl{[}q_{S_{0}}^{\ast}(\theta),B^{0,1}\bigr{]}=0$, we obtain (¯$(\overline{\partial}+\mathop{\rm ad}\nolimits\theta)(\tau)=\overline{\partial}_{E}(\widetilde{v}(B^{1,0}))+[\theta,\widetilde{v}(B^{0,1})]=0$.
There exists the following naturally defined quasi-isomorphisms:
@@
Let us compare the cocycles induced by $\tau$ and $\iota_{v}(\mathcal{E},\theta_{\mathcal{E}})$ in the complex $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr{)}$. In the following argument, we shall identify $\mathcal{E}$ and $q_{S_{0}}^{-1}(E)$ by $\varphi$, and we omit to denote $\varphi$. Let $X=\bigcup_{i=0}^{m}U_{i}$ be an open covering. We may assume that there exist holomorphic frames ${\boldsymbol{v}}_{i}$ of $\mathcal{E}_{|S_{0}\times U_{i}}$. We obtain the matrix valued holomorphic function $D_{i,j}$ on $S_{0}\times(U_{i}\cap U_{j})$ determined by ${\boldsymbol{v}}_{j|S_{0}\times(U_{i}\cap U_{j})}={\boldsymbol{v}}_{i|S_{0}\times(U_{i}\cap U_{j})}D_{ij}$. We also obtain the matrix valued holomorphic $1$-form $\Theta_{i}$ on $S_{0}\times U_{i}$ determined by $\theta_{\mathcal{E}}{\boldsymbol{v}}_{i}={\boldsymbol{v}}_{i}\Theta_{i}$. The restriction of ${\boldsymbol{v}}_{i}$ to $\{s_{0}\}\times U_{i}$ are denoted by ${\boldsymbol{v}}_{i|s_{0}}$. Let $\kappa_{i}$ be the holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes\Omega_{U_{i}}^{1}$ determined by $\kappa_{i}{\boldsymbol{v}}_{i|s_{0}}={\boldsymbol{v}}_{i|s_{0}}\widetilde{v}(\Theta_{i})$. Let $\rho_{i,j}$ be the holomorphic section of $\mathop{\rm End}\nolimits(E)_{U_{i}\cap U_{j}}$ determined by $\rho_{i,j}{\boldsymbol{v}}_{j}={\boldsymbol{v}}_{i}\widetilde{v}(D_{i,j})$. By definition, ($(\{\rho_{i,j}\},\{\kappa_{i}\})$ is the $1$-cocycle of $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta)\bigr{)}$ induced by $\iota_{v}(\mathcal{E})$. Let $M_{i}$ be the matrix valued $C^{\infty}$-function determined by ${\boldsymbol{v}}_{i}=q^{-1}_{S_{0}}({\boldsymbol{v}}_{i|s_{0}})M_{i}$. The entries of $M_{i}$ are contained in $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S_{0}}$. Because ¯$\overline{\partial}_{\mathcal{E}}(q_{S_{0}}^{-1}({\boldsymbol{v}}_{i|s_{0}}))=q_{S_{0}}^{-1}({\boldsymbol{v}}_{i|s_{0}})(-\overline{\partial}(M_{i})M_{i}^{-1})$, we obtain
@@
Hence, the $1$-cocycles induced by $\tau$ and $\iota_{v}(\mathcal{E})$ are cohomologue in the complex $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr{)}$.
@@
For a different $C^{\infty}$-isomorphism in Lemma 2.18, we obtain another $1$-cocycle $\tau^{\prime}$ of $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$. The difference $\tau-\tau^{\prime}$ is a $1$-coboundary. It follows from Lemma 2.19. We can also check it directly.
@@ 2.4.1 Harmonic $1$-forms of $E$
Let $h$ be a harmonic metric of $(E,\theta)$. Let =¯$\nabla=\overline{\partial}_{E}+\partial_{E,h}$ be the Chern connection of $(E,\theta)$. Let $\theta^{\dagger}_{h}$ denote the adjoint of $\theta$ with respect to $h$. We fix a conformal metric $g_{X}$ of $X$. Let (¯$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}:A^{i+1}(E)\to A^{i}(E)$ denote the formal adjoint of ¯$\overline{\partial}_{E}+\theta$ with respect to $h$ and $g_{X}$.
@@
$\tau\in A^{1}(E)$ is called a harmonic $1$-form of $(E,\theta,h)$ if (¯$(\overline{\partial}_{E}+\theta)\tau=0$ and (¯$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}\tau=0$. Here, (¯$(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}$ denotes adjoint of ¯$\overline{\partial}_{E}+\theta$ with respect to $h$ and $g_{X}$.
@@
Let $\Delta_{h,g_{X}}=(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}(\overline{\partial}_{E}+\theta)+(\overline{\partial}_{E}+\theta)(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}$. When $X$ is compact, we have $\Delta_{h,g_{X}}\tau=0$ if and only if (¯$(\overline{\partial}_{E}+\theta)\tau=(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}\tau=0$.
@@
$\tau$ is a harmonic $1$-form if and only if (¯$(\overline{\partial}_{E}+\theta)\tau=(\partial_{E,h}+\theta_{h}^{\dagger})\tau=0$.
Proof   Let $\Lambda_{X}$ denote the adjoint of the multiplication of the Kähler form associated with $g_{X}$. According to [25, Lemma 3.1], we have
(¯$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}\tau=-\sqrt{-1}[\Lambda_{X},\partial_{E,h}+\theta_{h}^{\dagger}]\tau=-\sqrt{-1}\Lambda_{X}((\partial_{E,h}+\theta_{h}^{\dagger})\tau).$
It is 00 if and only if $(\partial_{E,h}+\theta_{h}^{\dagger})\tau=0$.
@@
Let $s$ be a section of $E$ such that (¯$(\overline{\partial}_{E}+\theta)s=0$. Then, $\tau=(\partial_{E,h}+\theta_{h}^{\dagger})s$ is a harmonic $1$-form of $(E,\theta,h)$. Similarly, if $s^{\prime}$ is a section of $E$ such that $(\partial_{E,h}+\theta^{\dagger}_{h})s^{\prime}=0$, then (¯$(\overline{\partial}_{E}+\theta)s^{\prime}$ is a harmonic $1$-form of $(E,\theta,h)$.
@@
(¯$(\overline{\partial}_{E}+\theta)\tau=(\overline{\partial}_{E}+\theta)(\partial_{E,h}+\theta_{h}^{\dagger})s=-(\partial_{E,h}+\theta_{h}^{\dagger})(\overline{\partial}_{E}+\theta)s=0.$
@@
Suppose that there exists positive constants $B_{i}$ $(i=1,2)$ such that $|\tau_{2}|_{h,g_{X}}\leq B_{1}$ and $|\sigma|_{h}\leq B_{2}$ on a neighbourhood of $\partial U$. Then, there exists $B_{3}>0$, depending only on $B_{i}$ $(i=1,2)$ such that
$\|\tau_{|U}\|_{L^{2},h}\leq B_{3}.$
Here, $\|\tau_{|U}\|_{L^{2},h}$ denotes the $L^{2}$-norm of $\tau_{|U}$ with respect to $h$ and $g_{X}$. Note that it is independent of the choice of a conformal metric $g_{X}$.
@@ 2.4.2 Harmonic $1$-forms of $\mathop{\rm End}\nolimits(E)$
Let $X^{\dagger}$ denote the conjugate of $X$, and let $E^{\dagger}_{h}$ denote the holomorphic vector bundle $(E,\partial_{E,h})$ on $X^{\dagger}$. Then, $\theta^{\dagger}_{h}$ is a Higgs field of $E^{\dagger}_{h}$, and $h$ is a harmonic metric of $(E^{\dagger}_{h},\theta^{\dagger}_{h})$. Because $\partial_{E,h}(s^{\dagger}_{h})=(\overline{\partial}_{E}s)^{\dagger}_{h}=0$ and $[\theta^{\dagger}_{h},s^{\dagger}_{h}]=-[\theta,s]^{\dagger}_{h}=0$, $s^{\dagger}_{h}$ is a holomorphic section of $\mathcal{H}^{0}(E^{\dagger},\theta^{\dagger}_{h})$. Hence, (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta^{\dagger}_{h})s^{\dagger}_{h}$ is also a harmonic $1$-form of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$.
@@ 2.5.1 Uhlenbeck theorem
Let $V$ be a holomorphic vector bundle on $\mathcal{B}(1)$ equipped with a Hermitian metric $h$. Let $\nabla_{h}$ denote the Chern connection, and let $R(h)=\mathfrak{R}\,dz\,d\overline{z}$ denote the curvature. For a $C^{\infty}$-section $s$ of $V$ or $\mathop{\rm End}\nolimits(V)$ and for $10<r_{1}<$ 1, let $\|s_{|\mathcal{B}(r_{1})}\|_{L_{\ell}^{p},h,g_{0}}$ denote the $L_{\ell}^{p}$-norm of $s_{|\mathcal{B}(r_{1})}$ with respect to $\nabla_{h}$, $h$ and $g_{0}$. For a frame ${\boldsymbol{v}}$ of $V$, let $\mathcal{A}^{{\boldsymbol{v}}}$ denote the connection form of $\nabla_{h}$ with respect to ${\boldsymbol{v}}$. If $r_{1}=1$, $\|s_{|\mathcal{B}(1)}\|_{L_{\ell}^{p},h,g_{0}}$ is denoted by $\|s\|_{L_{\ell}^{p},h,g_{0}}$. We recall the following theorem of Uhlenbeck. (See [27, Theorem 1.3] for more precise statement.)
@@
The following lemma is also well known.
Lemma 2.33
Let $\ell\in{\mathbb{Z}}_{\geq 0}$ and $p>2$. Let $10<r_{1}<r_{2}<$ 1. There exists a positive constant $C_{1}(\ell,p,r_{1},r_{2})$ depending only on $\ell,p,r_{1},r_{2}$ such that the following holds.
  • Suppose that $|\mathfrak{R}|_{h}\leq\epsilon_{0}$ on $\mathcal{B}(1)$. Let ${\boldsymbol{v}}$ be a frame as in Theorem 2.32. Then,
    $\bigl{\|}\mathcal{A}^{{\boldsymbol{v}}}_{|\mathcal{B}(r_{1})}\bigr{\|}_{L_{\ell}^{p},g_{0}}\leq C_{1}(\ell,p,r_{1},r_{2})\|\mathfrak{R}_{|\mathcal{B}(r_{2})}\|_{L_{\ell}^{p},h,g_{0}}.$ (24)
Proof   Let $\mathcal{R}^{{\boldsymbol{v}}}$ be the matrix valued function determined by $\mathfrak{R}{\boldsymbol{v}}={\boldsymbol{v}}\cdot\mathcal{R}^{{\boldsymbol{v}}}$. The following holds.
$d(\mathcal{A}^{{\boldsymbol{v}}})+\mathcal{A}^{{\boldsymbol{v}}}\wedge\mathcal{A}^{{\boldsymbol{v}}}=\mathcal{R}^{{\boldsymbol{v}}}\,dz\,d\overline{z},\quad d^{\ast}(\mathcal{A}^{{\boldsymbol{v}}})=0$
We can obtain (24) by using the elliptic regularity and a standard bootstrapping argument.

### v3

@@ 2.3.3 Appendix: $C^{\infty}$-description of the induced infinitesimal deformation
For any complex submanifold $S^{\prime}$ of $S$, let $(\mathcal{E}_{S^{\prime}},\theta_{\mathcal{E},S^{\prime}})$ denote the restriction of $(\mathcal{E},\theta_{\mathcal{E}})$ to $S^{\prime}\times X$. Let $q_{S^{\prime}}:S^{\prime}\times X\to X$ denote the projection. Let $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S^{\prime}}$ denote the sheaf of $C^{\infty}$-functions $f$ on $S^{\prime}\times X$ such that $\overline{\partial}_{S^{\prime}}f=0$.
@@
Let $\varphi_{U_{0}}$ be the isomorphism $q_{S_{2}}^{\ast}(\mathcal{E})_{|S_{2}\times U_{0}}\simeq\mathcal{E}_{|S_{2}\times U_{0}}$ defined by $\varphi_{U_{0}}(q_{S_{2}}^{\ast}({\boldsymbol{v}}_{0|\{s_{0}\}\times U_{0}}))={\boldsymbol{v}}_{0|S_{2}\times U_{0}}$. Let $\varphi_{U_{i}}$ be the $C^{\infty}$-isomorphism $q_{S_{2}}^{\ast}(\mathcal{E})_{|S_{2}\times U_{i}}\simeq\mathcal{E}_{|S_{2}\times U_{i}}$ defined by
@@
Proof  Because $\overline{\partial}_{q_{S_{0}}^{\ast}(E)}(B^{1,0})+\bigl[q_{S_{0}}^{\ast}(\theta),B^{0,1}\bigr]=0$, we obtain $(\overline{\partial}+\mathop{\rm ad}\nolimits\theta)(\tau)=\overline{\partial}_{E}(\widetilde{v}(B^{1,0}))+[\theta,\widetilde{v}(B^{0,1})]=0$.
There exist the following naturally defined quasi-isomorphisms:
@@
Let us compare the cocycles induced by $\tau$ and $\iota_{v}(\mathcal{E},\theta_{\mathcal{E}})$ in the complex $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl(\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr)$. In the following argument, we shall identify $\mathcal{E}$ and $q_{S_{0}}^{-1}(E)$ by $\varphi$, and we omit to denote $\varphi$. Let $X=\bigcup_{i=0}^{m}U_{i}$ be an open covering. We may assume that there exist holomorphic frames ${\boldsymbol{v}}_{i}$ of $\mathcal{E}_{|S_{0}\times U_{i}}$. We obtain the matrix valued holomorphic function $D_{i,j}$ on $S_{0}\times(U_{i}\cap U_{j})$ determined by ${\boldsymbol{v}}_{j|S_{0}\times(U_{i}\cap U_{j})}={\boldsymbol{v}}_{i|S_{0}\times(U_{i}\cap U_{j})}D_{ij}$. We also obtain the matrix valued holomorphic $1$-form $\Theta_{i}$ on $S_{0}\times U_{i}$ determined by $\theta_{\mathcal{E}}{\boldsymbol{v}}_{i}={\boldsymbol{v}}_{i}\Theta_{i}$. The restriction of ${\boldsymbol{v}}_{i}$ to $\{s_{0}\}\times U_{i}$ are denoted by ${\boldsymbol{v}}_{i|s_{0}}$. Let $\kappa_{i}$ be the holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes\Omega_{U_{i}}^{1}$ determined by $\kappa_{i}{\boldsymbol{v}}_{i|s_{0}}={\boldsymbol{v}}_{i|s_{0}}\widetilde{v}(\Theta_{i})$. Let $\rho_{i,j}$ be the holomorphic section of $\mathop{\rm End}\nolimits(E)_{U_{i}\cap U_{j}}$ determined by $\rho_{i,j}{\boldsymbol{v}}_{j}={\boldsymbol{v}}_{i}\widetilde{v}(D_{i,j})$. By definition, $(\{\rho_{i,j}\},\{\kappa_{i}\})$ is the $1$-cocycle of $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl(\mathop{\rm Def}\nolimits(E,\theta)\bigr)$ induced by $\iota_{v}(\mathcal{E})$. Let $M_{i}$ be the matrix valued $C^{\infty}$-function determined by ${\boldsymbol{v}}_{i}=q^{-1}_{S_{0}}({\boldsymbol{v}}_{i|s_{0}})M_{i}$. The entries of $M_{i}$ are contained in $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S_{0}}$. Because $\overline{\partial}_{\mathcal{E}}(q_{S_{0}}^{-1}({\boldsymbol{v}}_{i|s_{0}}))=q_{S_{0}}^{-1}({\boldsymbol{v}}_{i|s_{0}})(-\overline{\partial}(M_{i})M_{i}^{-1})$, we obtain
@@
Hence, the $1$-cocycles induced by $\tau$ and $\iota_{v}(\mathcal{E})$ are cohomologous in the complex $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl(\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr)$.
@@
For a different $C^{\infty}$-isomorphism in Lemma 2.18, we obtain another 11-cocycle $\tau^{\prime}$ of $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$. The difference $\tau-\tau^{\prime}$ is a 11-coboundary. It follows from Lemma 2.19. We can also check it directly.
@@ 2.4.1 Harmonic $1$-forms of $E$
Let $h$ be a harmonic metric of $(E,\theta)$. Let $\nabla=\overline{\partial}_{E}+\partial_{E,h}$ be the Chern connection of $(E,\theta)$. Let $\theta^{\dagger}_{h}$ denote the adjoint of $\theta$ with respect to $h$. We fix a conformal metric $g_{X}$ of $X$. Let $(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}:A^{i+1}(E)\to A^{i}(E)$ denote the formal adjoint of $\overline{\partial}_{E}+\theta$ with respect to $h$ and $g_{X}$.
@@
$\tau\in A^{1}(E)$ is called a harmonic $1$-form of $(E,\theta,h)$ if $(\overline{\partial}_{E}+\theta)\tau=0$ and $(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}\tau=0$.
@@
Let $\Delta_{h,g_{X}}=(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}(\overline{\partial}_{E}+\theta)+(\overline{\partial}_{E}+\theta)(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}$. When $X$ is compact, we have $\Delta_{h,g_{X}}\tau=0$ if and only if $(\overline{\partial}_{E}+\theta)\tau=(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}\tau=0$.
@@
$\tau$ is a harmonic $1$-form if and only if $(\overline{\partial}_{E}+\theta)\tau=(\partial_{E,h}+\theta_{h}^{\dagger})\tau=0$.
Proof  Let $\Lambda_{X}$ denote the adjoint of the multiplication of the Kähler form associated with $g_{X}$. According to [26, Lemma 3.1], we have
$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}\tau=-\sqrt{-1}[\Lambda_{X},\partial_{E,h}+\theta_{h}^{\dagger}]\tau=-\sqrt{-1}\Lambda_{X}((\partial_{E,h}+\theta_{h}^{\dagger})\tau).$
It is 0 if and only if $(\partial_{E,h}+\theta_{h}^{\dagger})\tau=0$.
@@
Let $s$ be a section of $E$ such that $(\overline{\partial}_{E}+\theta)s=0$. Then, $\tau=(\partial_{E,h}+\theta_{h}^{\dagger})s$ is a harmonic $1$-form of $(E,\theta,h)$. Similarly, if $s^{\prime}$ is a section of $E$ such that $(\partial_{E,h}+\theta^{\dagger}_{h})s^{\prime}=0$, then $(\overline{\partial}_{E}+\theta)s^{\prime}$ is a harmonic $1$-form of $(E,\theta,h)$.
@@
$(\overline{\partial}_{E}+\theta)\tau=(\overline{\partial}_{E}+\theta)(\partial_{E,h}+\theta_{h}^{\dagger})s=-(\partial_{E,h}+\theta_{h}^{\dagger})(\overline{\partial}_{E}+\theta)s=0.$
@@
In Proposition 2.27, we moreover assume that $\tau_{1}=\tau_{2}$ and that there exist positive constants $B_{i}$ $(i=1,2)$ such that $|\tau_{1}|_{h,g_{X}}\leq B_{1}$ and $|\sigma_{1}|_{h}\leq B_{2}$ on a neighbourhood of $\partial U$. Then, there exists $B_{3}>0$, depending only on $B_{i}$ $(i=1,2)$ such that
$\|\tau_{1|U}\|_{L^{2},h}\leq B_{3}.$
Here, $\|\tau_{1|U}\|_{L^{2},h}$ denotes the $L^{2}$-norm of $\tau_{1|U}$ with respect to $h$ and $g_{X}$. Note that it is independent of the choice of a conformal metric $g_{X}$.
@@ 2.4.2 Harmonic $1$-forms of $\mathop{\rm End}\nolimits(E)$
Let $X^{\dagger}$ denote the conjugate of $X$, and let $E^{\dagger}_{h}$ denote the holomorphic vector bundle $(E,\partial_{E,h})$ on $X^{\dagger}$. Then, $\theta^{\dagger}_{h}$ is a Higgs field of $E^{\dagger}_{h}$, and $h$ is a harmonic metric of $(E^{\dagger}_{h},\theta^{\dagger}_{h})$. Because $\partial_{E,h}(s^{\dagger}_{h})=(\overline{\partial}_{E}s)^{\dagger}_{h}=0$ and $\theta^{\dagger}_{h},s^{\dagger}_{h}]=-[\theta,s]^{\dagger}_{h}=0$, $s^{\dagger}_{h}$ is a holomorphic section of $\mathcal{H}^{0}(E^{\dagger},\theta^{\dagger}_{h})$. Hence, $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta^{\dagger}_{h})s^{\dagger}_{h}$ is also a harmonic $1$-form of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$.
@@ 2.5.1 Uhlenbeck theorem
Let $V$ be a holomorphic vector bundle on $\mathcal{B}(1)$ equipped with a Hermitian metric $h$. Let $\nabla_{h}$ denote the Chern connection, and let $R(h)=\mathfrak{R}\,dz\,d\overline{z}$ denote the curvature. For a $C^{\infty}$-section $s$ of $V$ or $\mathop{\rm End}\nolimits(V)$ and for $0<r_{1}<1$, let $\|s_{|\mathcal{B}(r_{1})}\|_{L_{\ell}^{p},h,g_{0}}$ denote the $L_{\ell}^{p}$-norm of $s_{|\mathcal{B}(r_{1})}$ with respect to $\nabla_{h}$, $h$ and $g_{0}$. For a frame $\boldsymbol{v}$} of $V$, let $\mathcal{A}^{{\boldsymbol{v}}}$ denote the connection form of $\nabla_{h}$ with respect to $\boldsymbol{v}$}. If $r_{1}=1$, $\|s_{|\mathcal{B}(1)}\|_{L_{\ell}^{p},h,g_{0}}$ is denoted by $\|s\|_{L_{\ell}^{p},h,g_{0}}$. We recall the following theorem of Uhlenbeck. (See [28, Theorem 1.3] for more precise statement.)